OpenStudy (anonymous):

A drinking glass is shaped like a cylinder with a height of 12 cm and a radius of 4 cm. Maeva adds 25 spherical pieces of ice to the glass. The pieces of ice each have a diameter of 3 cm. How many cubic centimeters of water can Maeva add to fill the glass to the rim? Use 3.14 to approximate pi and express your answer in hundredths. cm3

OpenStudy (anonymous):

Do you know the formula for the volume of a cylinder?

OpenStudy (anonymous):

You there??

OpenStudy (anonymous):

Oh yeah im sorta busy multitasking

OpenStudy (anonymous):

sorry im really busy thats why i posted this question here

OpenStudy (anonymous):

i dont have much time but yeah i know the formula im not good at it though

OpenStudy (anonymous):

V=pi*r^2*h

OpenStudy (anonymous):

yep, but i gotta go its kinda too late, i hope i can get it sometime later but ill be back here on openstudy tomorrow before its required so ill probably see you tomorrow

OpenStudy (anonymous):

See ya then

OpenStudy (anonymous):

back can someone help

OpenStudy (mathstudent55):

You need to formula for the volume of a cylinder. Do you know it?

OpenStudy (anonymous):

duh V=pi*r^2*h someone just posted it

OpenStudy (anonymous):

but uh i suck at using it :c

OpenStudy (anonymous):

V=πr^2h = volume = 3.14x4^2x14

OpenStudy (anonymous):

703.36 for the cylinder

OpenStudy (anonymous):

14.14 each sphere multiplied by 25 = 353.5

OpenStudy (anonymous):

703.36-353.5=349.86 empty cubic centimeters that can be filled with water tell me if im right.

OpenStudy (mathstudent55):

The vol of the glass is: V = pi r^2 h = 3.14 * 4^2 * 12 = 602.88 cm^3

OpenStudy (mathstudent55):

The height of the glass is 12 cm, not 14 cm.

OpenStudy (mathstudent55):

Now you need the volume of the spheres of ice: The diameters of the spheres are 3 cm, so the radii are 1.5 cm. V = (4/3)pi r^3 V = (4/3) * 3.14 * (1.5)^3 14.13 cm^3 25 spheres have a volume of 25 * 14.13 = 353.25 cm^3 Now subtract 602.88 cm^3 - 353.25 cm^3

OpenStudy (anonymous):

mk so one small mistake thanks

OpenStudy (anonymous):

you forgot to medal me

OpenStudy (mathstudent55):

@Shawn1324 Great job! Your calculations were correct. The only little mistake was using the wrong number.

OpenStudy (anonymous):

yeah i got that mistake after, i made the mistake in the first place cause i was on the wrong question

OpenStudy (anonymous):

You got your part wrong by the way, i put an extra ? with a second best awnser with a different formula and the teach said that one was right but i didnt choose it. The one we did with your suggestion was wrong

OpenStudy (mathstudent55):

What did I do wrong? What is the correct answer?